Diffusion Morning Report 11/08/10

Ted Berg tberg at stanford.edu
Mon Nov 8 06:38:53 PST 2010


Good Morning All,
     Looking good this morning. Tylan 2 had MFS alarms during TLC clean. 
Epi had lift errors but has been run since. Tylan LTO  had MFS alarms. 
Ebeam has computer issues still. no other problems at this time.

					
	Date: 	11/8/10 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*yellow* 	*mfs alarms* 	** 	** 	*2*
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*up* 	** 	*.25/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*20.1/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*12.5/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*14.11/30* 	** 	**
Tylan 11-LTO 	*yellow* 	*O2 MFS alarms* 	*27.86/30* 	** 	*3*
Tylan 12-FGA 	*up* 	*
* 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*up* 	** 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*yellow* 	*testing* 	** 	** 	**
AMAT Epi 	*yellow* 	*lift issues* 	** 	** 	*1*
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*computer issues* 	** 	** 	*4*
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*A on line 708 PSI* 	*changeover working 
well* 	** 	**



-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20101108/c084596c/attachment.html>


More information about the diffusion mailing list